# A Communication Consistent Approach to Signal Temporal Logic Task Decomposition in Multi-Agent Systems

### v2

@@ 5.2 Handling conflicting conjunctions
Proof 5.7.
For a given $d\in\mathcal{I}_{F}$ consider the set $C\in\mathfrak{C}_{rs}^{m}(d)\subset\mathfrak{C}_{rs}(d)$ such that, by definition of $\mathfrak{C}_{rs}(d)$ in (25), we have $a^{d},b^{d}]\subseteq\bigcup_{l\in C}[a^{l},b^{l}]$. Then by equation Prop. 4 it is known that (42) ensures the existence of at least one index $l\in C$ such that $\mathcal{B}_{rs}^{d}\cap\mathcal{B}_{rs}^{l}\supseteq\{\bm{\xi}\}\neq\emptyset$. Therefore, a conflict as per Fact 2 is avoided for the index set C. Now consider another index set $C^{\prime}\neq C$ such that $C^{\prime}\in\mathfrak{C}_{rs}(d)\setminus\mathfrak{C}_{rs}^{m}(d)$. By definition of $\mathfrak{C}_{rs}^{m}(d)$, it also holds $C\subset C^{\prime}$. It is now sufficient to note that the satisfaction (42) for one index $l\in C\subset C^{\prime}$ disproves again the condition for conflicting conjunction as per Fact 2 over $C^{\prime}$. Since the result is valid for any $d\in\mathcal{I}_{F}$, any $C\in\mathfrak{C}_{rs}^{m}(d)$ and any $C^{\prime}\in\mathfrak{C}_{rs}(d)\setminus\mathfrak{C}_{rs}^{m}(d)$, we have proved that conflicts as per type 2 do not arise over $\phi_{rs}$.
Similarly to Prop. 5.4, the set $\mathfrak{C}^{m}_{rs}(d)$ represents the set of subsets in $\mathfrak{C}_{rs}(d)$ for a given $d\in\mathcal{I}_{F}$, that is “minimal”, in the sense that any set of indices $C\in\mathfrak{C}^{m}_{rs}$ does not contain any set $C^{\prime}\in\mathfrak{C}_{rs}(d)$ thus reducing the number of subsets of indices of tasks to be checked for conflicting conjunctions. The same remarks highlighted for Prop. 5.4 are also valid for Prop. 5.6. The last result of this section is presented next.
Proposition 5.8.
@@
Proof 5.9.
For a given $d\in\mathcal{I}_{F}$ consider the set $D\in\mathfrak{D}_{rs}^{M}(d)\subset\mathfrak{D}_{rs}(d)$ such that, by definition of $\mathfrak{D}_{rs}(d)$ in (26), we have $a^{d},b^{d}]\subseteq\bigcap_{l\in D}[a^{l},b^{l}]$. Then by equation Prop. 4 it is known that (43) ensures that $\mathcal{B}_{rs}^{d}\cap\bigcap_{l\in D}\mathcal{B}_{rs}^{l}\supseteq\{\bm{\xi}\}$. Therefore, a conflict as per Fact 3 is avoided over $D$. Now consider another set of indices $D^{\prime}\neq D$ such that $D^{\prime}\in\mathfrak{D}_{rs}(d)\setminus\mathfrak{D}_{rs}^{M}(d)$ and, by definition, it holds $D\supset D^{\prime}$. Then the condition (43) directly disproves the condition for conflicting conjunction as per Fact 3 since, it holds $\mathcal{B}_{rs}^{d}\cap\bigcap_{l\in D^{\prime}}\mathcal{B}_{rs}^{l}\supseteq\mathcal{B}_{rs}^{d}\cap\bigcap_{l\in D}\mathcal{B}_{rs}^{l}\supseteq\{\bm{\xi}\}$. Since the argument is valid for any $d\in\mathcal{I}_{F}$, any $D\in\mathfrak{D}_{rs}(d)$ and any $D^{\prime}\in\mathfrak{D}_{rs}(d)\setminus\mathfrak{D}_{rs}^{M}(d)$, we have proved that conflicts as per type 3 do not arise over $\phi_{rs}$.
One last definition provides a compact and uniform way to verify the absence of any type of conflicting conjunctions over $\phi_{rs}=\land_{k\in\mathcal{K}_{rs}}\varphi_{rs}^{k}$ as per Prop. 5.4, 5.6 and 5.8. Namely, the union of all the subset of indices that can cause conflicting conjunctions of tasks is defined as
[…]
@@
where $\Xi_{rs}$ is the set of auxiliary variables applied to impose the conditions (41), (42) and (43), while $y_{\mathfrak{Q}_{rs}}$ is the index mapping relating each set of indices $Q\in\mathfrak{Q}_{rs}$ with an auxiliary variable $\bm{\xi}_{rs}^{q}$ such that $q=y_{\mathfrak{Q}_{rs}}(Q)$. With these new definitions, the conjunctions $\wedge_{l\in Q}\varphi_{rs}^{l}$ for all $Q\in\mathfrak{Q}_{rs}$ represent potential conflicting conjunctions over the edge $(r,s)\in\mathcal{E}_{\pi}$. To avoid conflicts, the satisfaction of (41), (42) and (43) over $(r,s)\in\mathcal{E}_{\pi}$ can be compactly written as
@@
such that $q=y_{\mathfrak{Q}_{rs}}(Q)$. It now remains to clarify how conflicting conjunctions as per Fact 4-5 can be avoided. In this respect, Assumption. 1 ensures that cycles of tasks as per Fact 4-5 can not arise during decomposition of $\mathcal{G}_{\psi}$ into $\mathcal{G}_{\bar{\psi}}$. Indeed, since only edges in the communication graph $\mathcal{G}_{c}$ are applied for the decomposition and since $\mathcal{G}_{c}$ is acyclic, then, by construction, the new task graph $\mathcal{G}_{\bar{\psi}}$ must necessarily be acyclic, for which conflicts of types 4 and 5 do not arise. The reason why conditions (27)-(28) are excluded in our decomposition approach is due to the fact that their enforcement leads to non-convex constraints over the parameter vectors $\bm{\eta}_{rs}^{l}$. Moreover, the possibility of decentralizing the proposed decomposition, as presented in Section 6.2, is lost if this last type of constraint is introduced. When the acyclicity assumption on $\mathcal{G}_{c}$ can not be enforced, a solution to avoid conflicting conjunctions of type 4 and 5 was provided in our previous work [21] by restricting the sets $\mathcal{B}_{rs}$ to be hyper-rectangles instead of general polytopes.
5.3 Handling communication consistency
Concerning communication consistency of the parametric tasks $\varphi_{rs}^{l},\;\forall l\in\bar{\mathcal{K}}_{rs}$ over the edges $(r,s)\in\ \mathcal{E}_{\pi}$, the following result is provided
Proposition 5.10.
Consider an edge $(r,s)\in\mathcal{E}_{\pi}$ and the parametric task $\varphi_{rs}^{l}$ with $l\in\bar{\mathcal{K}}_{rs}$ as specified in (40). Furthermore, consider the truth set $\mathcal{B}_{rs}^{l}=\mathcal{P}(A_{rs}^{l},\bm{c}_{rs}^{l},\alpha_{rs}^{l}\bm{z}_{rs}^{l})$ and parameter vector $\bm{\eta}_{rs}^{l}$ associated with $\varphi_{rs}^{l}$. Let the generator set $V_{rs}^{l}=\nu(\mathcal{P}(A_{rs}^{l},\bm{0},\bm{z}_{rs}^{l}))$ as per Prop. 1 and the selection matrix $S$ such that $p_{i}=S\bm{x}_{i}\;\forall i\in\mathcal{V}$. Then $\varphi_{rs}^{l}$ is communication consistent if
@@ 5.3 Handling communication consistency
We prove that (47) implies (17). First let the polytope $\mathcal{L}_{rs}=\{S\bm{e}_{rs}=\bm{p}_{s}-\bm{p}_{r}\;|\;\bm{e}_{rs}\in\mathcal{B}_{rs}^{l}\}$ represent the projection of $\mathcal{B}_{rs}^{l}$ over the space of relative positions of agent $r$ and $s$. Moreover, let the distance function $f(S\bm{e}_{rs})=\|S\bm{e}_{rs}\|^{2}-r^{2}_{c},\forall S\bm{e}_{rs}\in\mathcal{L}_{rs}$. By Def. 2 and (4) the generators of $\mathcal{L}_{rs}$ are given by the vectors $\{SG^{l}_{k}\bm{\eta}_{rs}^{l}\}_{k=1}^{|V^{l}_{rs}|}$, where $\{G^{l}_{k}\bm{\eta}_{rs}^{l}\}_{k=1}^{|V^{l}_{rs}|}=\nu(\bar{\mathcal{B}}_{rs}^{l})=\nu(\mathcal{P}(A_{rs}^{l},\bm{c}_{rs}^{l},\alpha_{rs}^{l}\bm{z}_{rs}^{l}))$. Thus $S\bm{e}_{rs}=\sum_{k}\lambda_{k}SG^{l}_{k}\bm{\eta}_{rs}^{l},\;\forall S\bm{e}_{rs}\in\mathcal{L}_{rs}$ with $\lambda_{k}>0$, $\sum_{k}\lambda_{k}=1$. From, Jensen’s inequality [33, Thm. 3.4] we also have that $f(SG^{l}_{k}\bm{\eta}_{rs}^{l})\leq 0,\;\forall k=1,\ldots|V_{rs}^{l}|\Rightarrow f(S\bm{e}_{rs})=\|S\bm{e}_{rs}\|^{2}-r_{c}^{2}=\|\bm{p}_{r}-\bm{p}_{s}\|^{2}-r_{c}^{2}\leq 0,\;\forall S\bm{e}_{rs}\in\mathcal{L}_{rs}$. It is then sufficient to note that $f(SG^{l}_{k}\bm{\eta}_{rs}^{l})\leq 0,\;\forall k=1,\ldots|V_{rs}^{l}|$ is equivalent to (47) such that $\mathcal{B}_{rs}^{l}\cap\{\bm{e}_{rs}\in\mathbb{X}_{rs}\;|\;\|S\bm{e}_{rs}\|\leq r_{c}\}=\mathcal{L}_{rs}$, which implies (17).
@@
Note that $N^{l}_{k}$ are positive semi-definite matrices, and (47) is thus convex in $\bm{\eta}_{rs}^{l}$.
6 Parameters Optimization
In this section, the convex optimization to obtain optimal parameters $\bm{\eta}_{rs}^{l},\;\forall(r,s)\in\mathcal{E}_{\pi},\;\forall l\in\bar{\mathcal{K}}_{rs}$ is presented, with the constraints required for ensuring conflict-free conjunctions of tasks over the communication consistent edges $\mathcal{E}_{\pi}$ as well as communication consistency.
6.1 Centralized parameters optimization
6.1.1 Variables
@@
The variables of the optimization program are the parameters $\bm{\eta}_{rs}^{l},\;\forall(r,s)\in\mathcal{E}_{\pi}$ of the tasks $\varphi_{rs}^{l}$ introduced for the decomposition for all $l\in\bar{\mathcal{K}}_{rs}$ and all $(r,s)\in\mathcal{E}_{\pi}$, together with all the auxiliary variables $\bm{\xi}_{rs}^{q}\in\Xi_{rs},\;\forall(r,s)\in\mathcal{E}_{\pi},\;\forall q=1,\ldots|\Xi_{rs}|$ (as per (45a)) to avoid conflicts as explained in Sec. 5.2. We define the set of parameters $\mathcal{H}_{rs}$ for each edge $(r,s)$ as $\mathcal{H}_{rs}=\{\bm{\eta}_{rs}^{l}\,|\,l\in\bar{\mathcal{K}}_{rs}\}$. We recall that the mapping $y_{\Pi_{rs}}$ as per (38) assigns a unique index $l\in\bar{\mathcal{K}}_{rs}$ to each inconsistent edge $(i,j)\in\Pi_{rs}$ passing though $(r,s)$. Indeed, when an inconsistent task $\varphi_{ij}$ with decomposition path $\bm{\pi}_{i}^{j}$ passes through $(r,s)$, then a collaborative task $\bar{\varphi}_{rs}^{ij}\underset{y_{\Pi_{rs}}}{\equiv}\varphi_{rs}^{l}$, with $l\in\bar{\mathcal{K}}_{rs}$, is introduced over $(r,s)$.
6.1.2 Cost Function

### v3

@@ V-B Handling conflicting conjunctions during decomposition
Proof:
For a given $d\in\mathcal{I}_{F}$ consider the set $C\in\mathfrak{C}_{rs}^{m}(d)\subset\mathfrak{C}_{rs}(d)$ such that, by definition of $\mathfrak{C}_{rs}(d)$ in (27), we have $a^{d},b^{d}]\subseteq\bigcup_{l\in C}[a^{l},b^{l}]$. Then by equation Prop. 4 it is known that (43) ensures the existence of at least one index $l\in C$ such that $\mathcal{B}_{rs}^{d}\cap\mathcal{B}_{rs}^{l}\supseteq\{\bm{\xi}\}\neq\emptyset$. Therefore, a conflict as per Fact 2 is avoided for the index set C. Now consider another index set $C^{\prime}\neq C$ such that $C^{\prime}\in\mathfrak{C}_{rs}(d)\setminus\mathfrak{C}_{rs}^{m}(d)$. By definition of $\mathfrak{C}_{rs}^{m}(d)$, it also holds $C\subset C^{\prime}$. It is now sufficient to note that the satisfaction (43) for one index $l\in C\subset C^{\prime}$ disproves again the condition for conflicting conjunction as per Fact 2 over $C^{\prime}$. Since the result is valid for any $d\in\mathcal{I}_{F}$, any $C\in\mathfrak{C}_{rs}^{m}(d)$ and any $C^{\prime}\in\mathfrak{C}_{rs}(d)\setminus\mathfrak{C}_{rs}^{m}(d)$, we have proved that conflicts as per type 2 do not arise over $\phi_{rs}$. ∎
Similarly to Prop. 6, the set $\mathfrak{C}^{m}_{rs}(d)$ represents the set of subsets in $\mathfrak{C}_{rs}(d)$ for a given $d\in\mathcal{I}_{F}$, that is “minimal”, in the sense that any set of indices $C\in\mathfrak{C}^{m}_{rs}$ does not contain any set $C^{\prime}\in\mathfrak{C}_{rs}(d)$ thus reducing the number of subsets of indices of tasks to be checked for conflicting conjunctions. The same remarks highlighted for Prop. 6 are also valid for Prop. 7. The last result of this section is presented next.
Proposition 8
@@
Proof:
For a given $d\in\mathcal{I}_{F}$ consider the set $D\in\mathfrak{D}_{rs}^{M}(d)\subset\mathfrak{D}_{rs}(d)$ such that, by definition of $\mathfrak{D}_{rs}(d)$ in (26), we have $a^{d},b^{d}]\subseteq\bigcap_{l\in D}[a^{l},b^{l}]$. Then by equation Prop. 4 it is known that (44) ensures that $\mathcal{B}_{rs}^{d}\cap\bigcap_{l\in D}\mathcal{B}_{rs}^{l}\supseteq\{\bm{\xi}\}$. Therefore, a conflict as per Fact 3 is avoided over $D$. Now consider another set of indices $D^{\prime}\neq D$ such that $D^{\prime}\in\mathfrak{D}_{rs}(d)\setminus\mathfrak{D}_{rs}^{M}(d)$ and, by definition, it holds $D\supset D^{\prime}$. Then the condition (44) directly disproves the condition for conflicting conjunction as per Fact 3 since, it holds $\mathcal{B}_{rs}^{d}\cap\bigcap_{l\in D^{\prime}}\mathcal{B}_{rs}^{l}\supseteq\mathcal{B}_{rs}^{d}\cap\bigcap_{l\in D}\mathcal{B}_{rs}^{l}\supseteq\{\bm{\xi}\}$. Since the argument is valid for any $d\in\mathcal{I}_{F}$, any $D\in\mathfrak{D}_{rs}(d)$ and any $D^{\prime}\in\mathfrak{D}_{rs}(d)\setminus\mathfrak{D}_{rs}^{M}(d)$, we have proved that conflicts as per type 3 do not arise over $\phi_{rs}$. ∎
One last definition provides a compact and uniform way to verify the absence of any type of conflicting conjunctions over $\phi_{rs}=\land_{k\in\mathcal{K}_{rs}}\varphi_{rs}^{k}$ as per Prop. 6, 7 and 8. Namely, let
$\mathfrak{Q}_{rs}$ $=\bigl{(}\bigcup_{d\in\mathcal{I}_{F}}\{\;\{d\}\cup\{l\}\;|\;l\in C,\;\forall C\in\mathfrak{C}_{rs}^{m}(d)\}\bigr{)}\cup$ (45)
$\hskip 11.38092pt\bigl{(}\bigcup_{d\in\mathcal{I}_{F}}\{\;\{d\}\cup D\;|\;\forall D\in\mathfrak{D}_{rs}^{M}(d)\}\bigr{)}\cup\mathfrak{L}^{M}_{rs},$
@@
where $\Xi_{rs}$ is the set of auxiliary variables applied to impose the conditions (42), (43) and (44), while $y_{\mathfrak{Q}_{rs}}$ is the index mapping relating each set of indices $Q\in\mathfrak{Q}$ with an auxiliary variable $\bm{\xi}_{rs}^{q}$ such that $q=y_{\mathfrak{Q}_{rs}}(Q)$. With these new definitions, the conjunctions $\wedge_{l\in Q}\varphi_{rs}^{l}$ for all $Q\in\mathfrak{Q}_{rs}$ represent potential conflicting conjunctions over the edge $(r,s)\in\mathcal{E}_{\pi}$. To avoid conflicts, the satisfaction of (42), (43) and (44) over $(r,s)\in\mathcal{E}_{\pi}$ can be compactly written as
@@
such that $q=y_{\mathfrak{Q}_{rs}}(Q)$.
It now remains to clarify how conflicting conjunctions as per Fact 4-5 can be avoided. In this respect, Assumption. 1 ensures that cycles of tasks as per Fact 4-5 can not arise during decomposition of $\mathcal{G}_{\psi}$ into $\mathcal{G}_{\bar{\psi}}$. Indeed, since only edges in the communication graph $\mathcal{G}_{c}$ are applied for the decomposition and since $\mathcal{G}_{c}$ is acyclic, then, by construction, the new task graph $\mathcal{G}_{\bar{\psi}}\subseteq\mathcal{G}_{c}$ must necessarily be acyclic, for which conflicts of types 4 and 5 do not arise. The reason why conditions (28)-(29) are excluded in our decomposition approach is because their enforcement leads to non-convex constraints over the parameters $\bm{\eta}$ as briefly explained next. Consider, for consistency with the presentation of Fact 4-5, a conjunction of tasks $\land_{(r,s)\in\bm{\omega}}\varphi_{rs}$ for some cycle $\bm{\omega}$ and let the associated non-similar polytopic truth sets $\mathcal{B}_{rs}=\mathcal{P}(A_{rs},\bm{c}_{rs},\alpha_{rs},\bm{z}_{rs}),\;\forall(r,s)\in\epsilon(\bm{\omega})$ with $\bm{\eta}_{rs}=[\bm{c}_{rs},\alpha_{rs}]$ being parameters to be optimised. Then, let for brevity, $V_{rs}=\nu(\mathcal{B}_{rs})=\{\bm{v}_{rs}^{k}\}_{k=1}^{|V_{rs}|},\;\forall(r,s)\in\epsilon(\bm{\omega})$ where, by Prop. 2 we have that each vector $\bm{v}^{k}_{rs}(\bm{\eta}_{rs})\in V_{rs}$ is expressed as a linear function in $\bm{\eta}_{rs}$. Moreover, let $V^{\oplus}=\bigoplus_{(r,s)\in\epsilon(\bm{\omega})}V_{rs}=\{\bm{v}_{\oplus}^{k}(\bm{\eta}_{\oplus})\}_{k=1}^{|V^{\oplus}|}$, such that $\bm{\eta}_{\oplus}=[\bm{\eta}_{rs}]_{(r,s)\in\epsilon(\bm{\omega})}$, and such that each vector $\bm{v}_{\oplus}^{k}(\bm{\eta_{\oplus}})\in V^{\oplus}$ depends linearly on $\bm{\eta}_{\oplus}$. Noting that $\bigoplus_{(r,s)\in\epsilon(\bm{\omega})}\mathcal{B}_{rs}=conv(\bigoplus_{(r,s)\in\epsilon(\bm{\omega})}V_{rs})$ [38, Ch. 15, pp. 316], the definition of convex hull in Def. 2 can be applied to enforce the inclusion $\bm{0}\in\bigoplus_{(r,s)\in\epsilon(\bm{\omega})}\mathcal{B}_{rs}$ by introducing the coefficients $\lambda_{k}\geq 0,\;k=1,\ldots|V^{\oplus}|$ with $\sum_{k=1}^{|V^{\oplus}|}\lambda_{k}=1$ and such that $\sum_{k=1}^{|V^{\oplus}|}\lambda_{k}\bm{v}^{k}_{\oplus}(\bm{\eta}_{\oplus})=\bm{0}$. However, this last summation is a non-convex constraint in the variables $\lambda_{i},\forall k=1,\ldots|V^{\oplus}|$ and $\bm{\eta}_{\oplus}$. Moreover, the decentralization property of our decomposition approach, as presented in Section VI-B, is lost if this last type of constraint is introduced. When the acyclicity assumption on $\mathcal{G}_{c}$ can not be enforced, a solution to avoid conflicting conjunctions of type 4 and 5 was provided in our previous work [21] by restricting the sets $\mathcal{B}_{rs}$ to be hyper-rectangles instead of general polytopes.
V-C Handling communication consistency
Concerning communication consistency of the parametric tasks $\varphi_{rs}^{l},\;\forall l\in\bar{\mathcal{K}}_{rs}$ over the edges $(r,s)\in\ \mathcal{E}_{\pi}$, this can be guaranteed by imposing appropriate constraints over the parameter vectors $\bm{\eta}_{rs}^{l},\;\forall l\in\bar{\mathcal{K}}_{rs}$ according to the following result.
Proposition 9
Consider an edge $(r,s)\in\mathcal{E}_{\pi}$ and the parametric task $\varphi_{rs}^{l}$ with $l\in\bar{\mathcal{K}}_{rs}$ as specified in (41). Furthermore, consider the truth set $\mathcal{B}_{rs}^{l}=\mathcal{P}(A_{rs}^{l},\bm{c}_{rs}^{l},\alpha_{rs}^{l}\bm{z}_{rs}^{l})$ and parameter vector $\bm{\eta}_{rs}^{l}$ associated with $\varphi_{rs}^{l}$. Let the generator set $V_{rs}^{l}=\nu(\mathcal{P}(A_{rs}^{l},\bm{0},\bm{z}_{rs}^{l}))$ as per Prop. 1. Then $\varphi_{rs}^{l}$ is communication consistent if
@@ V-C Handling communication consistency
We prove that (48) implies (18). First let the polytope $\mathcal{L}_{rs}=\{S\bm{e}_{rs}=\bm{p}_{s}-\bm{p}_{r}\;|\;\bm{e}_{rs}\in\mathcal{B}_{rs}^{l}\}$ represent the projection of $\mathcal{B}_{rs}^{l}$ over the space of relative positions of agent $r$ and $s$. Moreover, let the distance function $f(S\bm{e}_{rs})=\|S\bm{e}_{rs}\|^{2}-r^{2}_{c},\forall S\bm{e}_{rs}\in\mathcal{L}_{rs}$. By Def. 2 and (5) the generators of $\mathcal{L}_{rs}$ are given by the vectors $\{SG^{l}_{k}\bm{\eta}_{rs}^{l}\}_{k=1}^{|V^{l}_{rs}|}$, where $\{G^{l}_{k}\bm{\eta}_{rs}^{l}\}_{k=1}^{|V^{l}_{rs}|}=\nu(\bar{\mathcal{B}}_{rs}^{l})=\nu(\mathcal{P}(A_{rs}^{l},\bm{c}_{rs}^{l},\alpha_{rs}^{l}\bm{z}_{rs}^{l}))$. Thus $S\bm{e}_{rs}=\sum_{k}\lambda_{k}SG^{l}_{k}\bm{\eta}_{rs}^{l},\;\forall S\bm{e}_{rs}\in\mathcal{L}_{rs}$ with $\lambda_{k}>0$, $\sum_{k}\lambda_{k}=1$. From, Jensen’s inequality [33, Thm. 3.4] we also have that $f(SG^{l}_{k}\bm{\eta}_{rs}^{l})\leq 0,\;\forall k=1,\ldots|V_{rs}^{l}|\Rightarrow f(S\bm{e}_{rs})=\|S\bm{e}_{rs}\|^{2}-r_{c}^{2}=\|\bm{p}_{r}-\bm{p}_{s}\|^{2}-r_{c}^{2}\leq 0,\;\forall S\bm{e}_{rs}\in\mathcal{L}_{rs}$. It is then sufficient to note that $f(SG^{l}_{k}\bm{\eta}_{rs}^{l})\leq 0,\;\forall k=1,\ldots|V_{rs}^{l}|$ is equivalent to (48) such that $\mathcal{B}_{rs}^{l}\cap\{\bm{e}_{rs}\in\mathbb{X}_{rs}\;|\;\|S\bm{e}_{rs}\|\leq r_{c}\}=\mathcal{L}_{rs}$, which implies (18). ∎
@@
Note that the matrices $N^{l}_{k}$ are positive semi-definite such that relations (48) are convex in $\bm{\eta}_{rs}^{l}$.
VI Parameters optimization
In Sec. V we have developed a framework to decompose communication inconsistent collaborative tasks in the form $\phi_{ij}=\varphi_{ij}$ by introducing parametric tasks within the family of parametric predicates in (31). Furthermore, we have provided relevant constraints on the parameters of the parametric tasks to ensure conflict-free conjunctions of tasks over the communication consistent edges $\mathcal{E}_{\pi}$ applied for the decomposition and to ensure communication consistency of these. In this section, we show how convex optimization can be leveraged to obtain optimal parameters $\bm{\eta}_{rs}^{l},\;\forall(r,s)\in\mathcal{E}_{\pi},\;\forall l\in\bar{\mathcal{K}}_{rs}$ that satisfy the aforementioned constraints and such that the decomposition accuracy as (36) is maximised.
VI-A Centralized parameters optimization
The convex optimization program to solve the decomposition of the inconsistent tasks is presented next by defining the variables, cost function and constraints.
VI-A1 Variables
@@
The variables of the optimization program are the parameters $\bm{\eta}_{rs}^{l},\;\forall(r,s)\in\mathcal{E}_{\pi}$ of the tasks $\varphi_{rs}^{l}$ introduces by the decomposition for all $l\in\bar{\mathcal{K}}_{rs}$ and all $(r,s)\in\mathcal{E}_{\pi}$, together with all the auxiliary variables $\bm{\xi}_{rs}^{q}\in\Xi_{rs},\;\forall(r,s)\in\mathcal{E}_{\pi},\;\forall q=1,\ldots|\Xi_{rs}|$ (as per (46a)) applied to avoid conflicting conjunctions as explained in Sec. V-C. We thus define the set of parameters $\mathcal{H}_{rs}$ for each edge $(r,s)$ as $\mathcal{H}_{rs}=\{\bm{\eta}_{rs}^{l}\,|\,l\in\bar{\mathcal{K}}_{rs}\}$. We recall that the mapping $y_{\Pi_{rs}}$ as per (39) assigns a unique index $l\in\bar{\mathcal{K}}_{rs}$ to each inconsistent edge $(i,j)\in\Pi_{rs}$ passing though $(r,s)$. Indeed, when an inconsistent task $\varphi_{ij}$ with decomposition path $\bm{\pi}_{i}^{j}$ passes through $(r,s)$, then a collaborative task $\bar{\varphi}_{rs}^{ij}\underset{y_{\Pi_{rs}}}{=}\varphi_{rs}^{l}$, with $l\in\bar{\mathcal{K}}_{rs}$, is introduced over $(r,s)$.
VI-A2 Cost Function